# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v1

@@ Abstract
The denoising diffusion probabilistic model (DDPM) has emerged as a mainstream generative model in generative AI. While sharp convergence guarantees have been established for the DDPM, the iteration complexity is, in general, proportional to the ambient data dimension, resulting in overly conservative theory that fails to explain its practical efficiency. This has motivated the recent work Li and Yan, 2024a to investigate how the DDPM can achieve sampling speed-ups through automatic exploitation of intrinsic low dimensionality of data. We strengthen this prior work by demonstrating, in some sense, optimal adaptivity to unknown low dimensionality. For a broad class of data distributions with intrinsic dimension $k$, we prove that the iteration complexity of the DDPM scales nearly linearly with $k$, which is optimal when using KL divergence to measure distributional discrepancy. Our theory is established based on a key observation: the DDPM update rule is equivalent to running a suitably parameterized SDE upon discretization, where the nonlinear component of the drift term is intrinsically low-dimensional.
@@ 1 Introduction
In generative modeling, one is asked to produce novel samples that resemble, in distribution, the data on which the model has been trained. In order to fully unleash its power, a generative modeling algorithm must exploit — either implicitly or explicitly — the distinctive features of the data distribution of interest, rather than merely optimizing for worst-case scenarios. A prominent example of such distinctive features is the low-dimensional structure underlying the data of interest, which permeates modern statistical modeling. For instance, image data often resides on some low-dimensional manifold (Simoncelli and Olshausen,, 2001; Pope et al.,, 2021), and genomics data might also be effectively modeled through manifolds (Moon et al.,, 2018; Zhu et al.,, 2018). While harnessing these low-dimensional data structures is generally believed to enable more efficient data generation, the theoretical underpinnings for this capability in contemporary generative modeling algorithms remain significantly under-explored.
In this paper, we study how low-dimensional data structure influences the performance of diffusion models, an emerging paradigm that facilitates remarkable progress across various subfields of modern generative modeling (Chen et al., 2024a, ; Sohl-Dickstein et al.,, 2015; Dhariwal and Nichol,, 2021; Ho et al.,, 2022; Ramesh et al.,, 2022; Watson et al.,, 2023). While empirical evidence suggests that diffusion models might not suffer from the curse of dimensionality, a rigorious understanding of their efficacy in the presence of low-dimensional structure remains elusive. Consider, for concreteness, a mainstream diffusion model known as the denoising diffusion probabilistic model (DDPM) (Ho et al.,, 2020), which has received widespread adoption in practice and sparked considerable interest from the theoretical community to demystify its efficacy (e.g., Chen et al., 2022b ; Chen et al., 2022a ; Lee et al., (2022, 2023); Li et al., 2024b ; Tang, (2023); Mbacke and Rivasplata, (2023); Tang and Zhao, (2024); Liang et al., (2024); Li and Yan, 2024b ). In order for the DDPM sampler to generate a $d$-dimensional new sample that approximates well the target distribution in terms of KL divergence, the number of steps [^1] Given that each step of the DDPM requires evaluating a score function (typically through computing the output of a large neural network or transformer), the sampling speed of the DDPM is largely dictated by the number of steps. needs to scale linearly with $d$, a scaling that has been shown to be un-improvable in general (Benton et al.,, 2024). However, in application like high-resolution image generation, the number of pixels in each image might be exceedingly large, far exceeding the number of steps typically needed to run the DDPM in practice. This indicates that even linear scaling with $d$ does not adequately explain the observed efficiency of the DDPM in practice.
@@
Despite this remarkable theoretical progress, however, it remains unclear to what extent the low-dimensional structure can enhance the sampling efficiency of the DDPM. More precisely, Li and Yan, 2024a proved that when the support of $p_{\mathsf{data}}$ has intrinsic dimension $k$ (to be made precise in Section 4.1.1), the iteration complexity of the DDPM scales at most with $k^{4}$ (up to some logarithmic factor and other dependency on the target accuracy level), provided that KL divergence is used to measure the distributional discrepancy. This result was recently improved to $k^{3}$ by Azangulov et al., (2024) under some low-dimensional smooth manifold assumption. This raises a natural question: is the $k^{3}$ scaling essential, and if not, how to sharpen it? Addressing this question would provide a quantitative understanding about DDPM’s degree of adaptivity when tackling low-dimensional structure.
@@
In this work, we develop a sharp convergence theory for the DDPM in the face of unknown low dimensionality. For a broad family of target data distributions $p_{\mathsf{data}}$ with intrinsic dimension $k$, we show that it takes DDPM at most $O(k/\varepsilon^{2})$ (up to some logarithmic factor) steps to yield an $\varepsilon^{2}$-accurate distribution — measured by the KL divergence — assuming access to perfect score functions. This improves upon the state-of-the-art iteration complexity (Azangulov et al.,, 2024) by a factor of $k^{2}$, and achieves nearly optimal scaling in $k$ without any burn-in requirement. Further, our convergence theory can be easily extended to accommodate imperfect score estimation.
Our convergence theory is established based on a key observation: the original DDPM update rule by Ho et al., (2020) is precisely equivalent to running a carefully parameterized stochastic differential equation (SDE) upon discretization using the exponential integrator. Crucially, the drift term of this discretized SDE is semi-linear, with the non-linear component proportional to the posterior mean of the data given its Gaussian corruption. Intuitively, when the support of $p_{\mathsf{data}}$ is low-dimensional, this nonlinear drift term acts as a “projection” onto the low-dimensional manifold of interest, harnessing this intrinsic structure to enhance the smoothness of the solution path and thereby speed up the sampling process. Somewhat surprisingly, while DDPM (Ho et al.,, 2020) was initially proposed to maximize a variational lower bound on the log-likelihood, the resulting update rule inherently adopts the desirable time discretization schedule, in a way that is fully adaptive to unknown low-dimensional distributions. A key ingredient underlying our analysis of the discretized SDE lies in carefully calibrating the evolution of the posterior mean function along the forward process.
While finalizing this manuscript, we became aware of an interesting concurrent work Potaptchik et al., (2024) that develops a similar convergence theory as ours. We shall discuss the difference between our work and Potaptchik et al., (2024) in Section 4.2.
@@ 2 Preliminaries
3 Key observation: DDPM as an adaptively discretized SDE
While previous studies (e.g., Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024)) have explored the convergence properties of time discretization of the ideal reverse process (5) through the SDE framework, these results fall short of directly addressing the precise update rule (2) of the DDPM sampler. As pointed out by Li and Yan, 2024a , the coefficients (or weights) adopted in the DDPM-type sampler are critical for effective adaptation to unknown low dimensionality; in fact, even mild perturbation of the coefficients can lead to considerable performance degradation, a curious phenomenon that does not occur in the general full-dimensional setting.
In order to rigorously pin down the effectiveness of the DDPM update rule (2), we first uncover its equivalence to a carefully constructed SDE. To be precise, defining
@@ 3 Key observation: DDPM as an adaptively discretized SDE
for the sake of notational convenience, we introduce the following SDE:
@@
    $\mathrm{d}Y_{t}$ $=\left(Y_{t}+2\left(\frac{\eta_{T-t}}{\eta_{T-t_{n}}}\cdot\frac{1}{\sigma^{2}_{T-t_{n}}}Y_{t_{n}}-\frac{1}{\sigma^{2}_{T-t}}Y_{t}\right)+\frac{2\eta_{T-t}}{\eta_{T-t_{n}}}\widehat{s}_{T-t_{n}}(Y_{t_{n}})\right)\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}$ (11a)
@@
With this result in place, the equivalence of the solution (12) and the original DDPM update rule (2) is self-evident. The proof of Theorem 1 is postponed to Section 5.1.
Recall that in the general full-dimensional settings, previous work (e.g., Chen et al., 2023a ) suggested that discretizing different parametrization of (5) typically results in the same order of discretization errors, indicating that distinguishing between them may be unnecessary. This message, however, does not hold in the presence of low-dimensional structure, which merits further discussions as follows.
@@
  • Parameterization matters. It is worth noting that how we re-parameterize the drift term of (5) — namely, how we rearrange the semi-linear drift term therein — can exert significant influences upon the sampling efficiency. This can be elucidated via the inspiring example singled out in Li and Yan, 2024a (, Theorem 2). More concretely, when the target data distribution is isotropic Gaussian on a $k$-dimension linear subspace ($k\ll d$), the iteration complexity of the DDPM sampler is proportional to $O(\mathrm{poly}(k))$ (Li and Yan, 2024a, ), but most of other DDPM-type update rules (i.e., the ones with different coefficients than the original DDPM sampler) can take $\Omega(\mathrm{poly}(d))$ steps to yield the target sampling quality. Informally speaking, parameterizing the semi-linear drift term differently might result in a solution within $O(d/\mathrm{poly}(N))$ KL divergence from the DDPM solution [^3] Roughly speaking, in view of the Taylor expansion, we know that the discrepancy between different solutions is of order $O(1/\mathrm{poly}(N))$; but since $d$-dimensional Gaussian noise is added in each step of the sampling procedure, this discrepancy can result in $O(d/\mathrm{poly}(N))$ difference in KL divergence per step.; this discrepancy is negligible in the full-dimensional setting, but could blow up when $k$ is exceedingly small compared to $d$ and when $N$ is proportional to $k$. All this underscores the critical importance of identifying a parameterization that adapts naturally to the intrinsic data geometry.
@@ 4.1.1 Low dimensionality of the target distribution
    and $\mathsf{dist}(x,\mathcal{M})$ denotes the Euclidean distance of $x$ to $\mathcal{M}$. Also, $\mathrm{vol}(\mathcal{M})$ is the volume, or equivalently, the $k$-dimensional Hausdorff measure, of the manifold $\mathcal{M}$.
    Roughly speaking, the reach of a manifold measures the scale of which the manifold can be regarded as a linear subspace. The larger the reach $\tau_{\mathcal{M}}$ is, the more regular the manifold appears. The volume of the manifold $\mathcal{M}$ is a measure of the size of the manifold, which generalizes the concept of volume in 3-dimensional Euclidean space. Instead of requiring data to be strictly supported on the manifold, Assumption 1 also allows the data to lie in the $\varepsilon_{0}$-enlargement of the manifold $\mathcal{M}$ defined as $\mathcal{M}^{\varepsilon_{0}}:=\{x\in\mathbb{R}^{d}\mid\mathsf{dist}(x,\mathcal{M})\leq\varepsilon_{0}\}$. This is stated formally in the following lemma, with the proof deferred to Appendix B.1.
@@
    If the support of the target data distribution $p_{\mathsf{data}}$ is contained within the manifold $\mathcal{M}^{\varepsilon_{0}}$, i.e., $\mathcal{X}_{\mathsf{data}}\subseteq\mathcal{M}^{\varepsilon_{0}}$,and if $\tau_{\mathcal{M}}=\Omega(k^{-C_{1}})$ and $\mathrm{vol}(\mathcal{M})=O(k^{C_{2}})$ for some universal constants $C_{1},C_{2}>0$, then Assumption 1 holds.
@@
  • Union of the above sets. Our framework can also cover a union of the structures mentioned above, that is, Assumption 1 automatically holds when the data are supported on a union of $k^{O(k)}$ subsets, each of which satisfies Assumption 1.
@@ 4.1.2 Discretization time points
Recall that the discretization time points are $0=t_{0}<t_{1}<\cdots<t_{N}<t_{N+1}=T$, with $T$ the time horizon of the forward process. In addition, we also define throughout the associated time intervals as
$\Delta_{n}\coloneqq t_{n+1}-t_{n}\quad(0\leq n\leq N)\qquad\text{and}\qquad\delta\coloneqq\Delta_{N}=T-t_{N}<1.$ (14)
In particular, $\delta$ shall often be referred to as the early stopping time, the meaning of which shall be discussed momentarily when presenting our convergence theory.
@@
$\Delta_{n}\leq\kappa\min\{1,T-t_{n}\}\qquad\text{for all }0\leq n\leq N.$ (15)
@@ 4.1.3 Score estimation errors
$\sum_{n=0}^{N-1}\Delta_{n}\mathbb{E}\left[\big{\|}s_{T-t_{n}}(X_{T-t_{n}})-\widehat{s}_{T-t_{n}}(X_{T-t_{n}})\big{\|}_{2}^{2}\right]\leq\varepsilon_{\mathsf{score}}^{2}<\infty,$
@@
Treating the score learning phase as a black box, Assumption 3 singles out a single quantity, $\varepsilon_{\mathsf{score}}\geq 0$, that captures the accuracy of the pre-trained score functions, which is commonly used in recent development of diffusion model theory (Chen et al., 2023a, ; Li et al.,, 2023; Li et al., 2024c, ; Benton et al.,, 2024; Dou et al.,, 2024; Li and Yan, 2024b, ; Li et al., 2024a, ; Li and Yan, 2024a, ; Li and Jiao,, 2024; Liang et al.,, 2024). Notably, assuming a weighted average of the mean squared score errors is clearly less stringent than requiring the mean squared score estimation errors to hold uniformly over all time discretization points (Chen et al., 2022b, ; Lee et al.,, 2022; Chen et al., 2023c, ; Wu et al.,, 2024).

### v2

@@ Abstract
The denoising diffusion probabilistic model (DDPM) has emerged as a mainstream generative model in generative AI. While sharp convergence guarantees have been established for the DDPM, the iteration complexity is, in general, proportional to the ambient data dimension, resulting in overly conservative theory that fails to explain its practical efficiency. This has motivated the recent work Li and Yan, 2024a to investigate how the DDPM can achieve sampling speed-ups through automatic exploitation of intrinsic low dimensionality of data. We strengthen this line of work by demonstrating, in some sense, optimal adaptivity to unknown low dimensionality. For a broad class of data distributions with intrinsic dimension $k$, we prove that the iteration complexity of the DDPM scales nearly linearly with $k$, which is optimal when using KL divergence to measure distributional discrepancy. Notably, our work is closely aligned with the independent concurrent work Potaptchik et al., (2024) — posted two weeks prior to ours — in establishing nearly linear-$k$ convergence guarantees for the DDPM.
@@ 1 Introduction
In generative modeling, one is asked to produce novel samples that resemble, in distribution, the data on which the model has been trained. In order to fully unleash its power, a generative modeling algorithm must exploit, either implicitly or explicitly, the distinctive features of the data distribution of interest, rather than merely optimizing for worst-case scenarios. A prominent example of such distinctive features is the low-dimensional structure underlying the data of interest, which permeates modern statistical modeling. For instance, image data often resides on some low-dimensional manifold (Simoncelli and Olshausen,, 2001; Pope et al.,, 2021), and genomics data might also be effectively modeled through manifolds (Moon et al.,, 2018; Zhu et al.,, 2018). While harnessing these low-dimensional data structures is generally believed to enable more efficient data generation, the theoretical underpinnings for this capability in contemporary generative modeling algorithms remain vastly under-explored.
In this paper, we study how low-dimensional data structure influences the performance of diffusion models, an emerging paradigm that facilitates remarkable progress across various subfields of modern generative modeling (Chen et al., 2024a, ; Sohl-Dickstein et al.,, 2015; Dhariwal and Nichol,, 2021; Ho et al.,, 2022; Ramesh et al.,, 2022; Watson et al.,, 2023). While empirical evidence suggests that diffusion models might not suffer from the curse of dimensionality, a rigorious understanding of their efficacy in the presence of low-dimensional structure remains elusive. Consider, for concreteness, a mainstream diffusion model known as the denoising diffusion probabilistic model (DDPM) (Ho et al.,, 2020), which has received widespread adoption in practice and sparked considerable interest from the theoretical community to demystify its efficacy (e.g., Chen et al., 2022b ; Chen et al., 2022a ; Lee et al., (2022, 2023); Li et al., 2024b ; Tang, (2023); Mbacke and Rivasplata, (2023); Tang and Zhao, (2024); Liang et al., (2024); Li and Yan, 2024b ). In order for the DDPM sampler to generate a $d$-dimensional new sample that approximates well the target distribution in terms of KL divergence, the number of steps [^1] Given that each step of the DDPM requires evaluating a score function (typically through computing the output of a large neural network or transformer), the sampling speed of the DDPM is largely dictated by the number of steps. needs to scale linearly with $d$, a scaling that has been shown to be un-improvable in general (Benton et al.,, 2024). However, in application like image generation, the number of pixels in each image might oftentimes be exceedingly large, far exceeding the number of steps typically needed to run the DDPM in practice. This indicates that even linear scaling with $d$ does not adequately explain the observed efficiency of the DDPM in practice.
@@
Despite this remarkable theoretical progress, however, it was unclear to what extent the low-dimensional structure could enhance the sampling efficiency of the DDPM. More precisely, Li and Yan, 2024a proved that when the support of $p_{\mathsf{data}}$ has intrinsic dimension $k$ (to be made precise in Section 4.1.1), the iteration complexity of the DDPM scales at most with $k^{4}$ (up to some logarithmic factor and other dependency on the target accuracy level), provided that KL divergence is used to measure the distributional discrepancy. This result was recently improved by Azangulov et al., (2024) to the order of $k^{3}$, assuming that data lies within some low-dimensional smooth manifolds. This raises a natural question: is the $k^{3}$ scaling essential, and if not, how to sharpen it? Addressing this question would provide a quantitative understanding about DDPM’s degree of adaptivity when tackling low-dimensional structure.
@@
In this work, we develop a sharp convergence theory for the DDPM in the face of unknown low dimensionality. For a broad family of target data distributions $p_{\mathsf{data}}$ with intrinsic dimension $k$, we show that it takes DDPM at most $O(k/\varepsilon^{2})$ (up to some logarithmic factor) steps to yield an $\varepsilon^{2}$-accurate distribution — measured by the KL divergence — assuming access to perfect score functions. This result, which shares similarity with a concurrent work Potaptchik et al., (2024) to be discussed momentarily, improves upon the iteration complexity derived by Azangulov et al., (2024) by a factor of $k^{2}$, and achieves nearly optimal scaling in $k$ without any burn-in requirement. Further, our convergence theory can be easily extended to accommodate imperfect score estimation.
To provide some intuition about its adaptivity, it is helpful to recognize that the original DDPM update rule by Ho et al., (2020) is precisely equivalent to running a carefully parameterized stochastic differential equation (SDE) upon discretization using the exponential integrator, a key observation that appeared in the recent work Azangulov et al., (2024). Crucially, the drift term of this discretized SDE is semi-linear, with the non-linear component proportional to the posterior mean of the data given its Gaussian corruption. Intuitively, when the support of $p_{\mathsf{data}}$ is low-dimensional, this nonlinear drift term acts as a “projection” onto the low-dimensional manifold of interest, harnessing this intrinsic structure to enhance the smoothness of the solution path and thereby speed up the sampling process. Somewhat surprisingly, while DDPM (Ho et al.,, 2020) was initially proposed to maximize a variational lower bound on the log-likelihood, the resulting update rule inherently adopts the desirable time discretization schedule, in a way that is fully adaptive to unknown low-dimensional distributions. A key ingredient underlying our analysis of the discretized SDE lies in carefully calibrating the evolution of the posterior mean function along the forward process.
The concurrent work Potaptchik et al., (2024).
While finalizing this work, we became aware of an interesting concurrent work Potaptchik et al., (2024), posted two weeks before our manuscript, that delivered the first convergence guarantees exhibiting linear dependency on the intrinsic manifold dimension of the data. Our paper, which stands as an independent and concurrent contribution, has similar aim and establishes similar results as Potaptchik et al., (2024), albeit under different assumptions on the low-dimensional structure. We shall provide more detailed comparisons between our work and Potaptchik et al., (2024) in Section 4.2.
@@ 2 Preliminaries
3 DDPM as an adaptively discretized SDE
While previous studies (e.g., Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024)) have explored the convergence properties of time discretization of the ideal reverse process (5) through the SDE framework, a dominant fraction of these results fell short of directly addressing the precise update rule (2) of the DDPM sampler. As pointed out by Li and Yan, 2024a , the coefficients (or weights) adopted in the DDPM-type sampler are critical for effective adaptation to unknown low dimensionality; in fact, even mild perturbation of the coefficients can lead to considerable performance degradation, a curious phenomenon that does not occur in the general full-dimensional setting.
In order to rigorously pin down the effectiveness of the DDPM update rule (2), a key step is to recognize its equivalence to a carefully constructed SDE. To be precise, set
@@ 3 DDPM as an adaptively discretized SDE
for notational convenience, and let us introduce the following SDE (11) that is provably equivalent to the DDPM update rule.
@@
    $\mathrm{d}Y_{t}$ $=\left(Y_{t}+2\left(\frac{\eta_{T-t}}{\eta_{T-t_{n}}}\cdot\frac{1}{\sigma^{2}_{T-t_{n}}}Y_{t_{n}}-\frac{1}{\sigma^{2}_{T-t}}Y_{t}\right)+\frac{2\eta_{T-t}}{\eta_{T-t_{n}}}\widehat{s}_{T-t_{n}}(Y_{t_{n}})\right)\mathrm{d}t+\sqrt{2}\,\mathrm{d}\mathcal{B}_{t}$ (11a)
@@
Remark 1.
After the initial posting of this paper, the authors of the concurrent work Potaptchik et al., (2024) informed us that a similar SDE already appeared in the recent work Azangulov et al., (2024, Section 7.2) as a type of backward SDE discretization with first-order score correction, where they discussed the motivation of score correction and its effect upon the discretization error. See also Azangulov et al., (2024, Page 90) for their justification of the SDE equivalence of the DDPM.
@@
With this result in place, the equivalence of the solution (12) and the original DDPM update rule (2) is self-evident. The proof of Theorem 1 is postponed to Appendix A.
Recall that in the general full-dimensional settings, previous work (e.g., Chen et al., 2023a ) suggested that discretizing different parametrization of (5) typically results in the same order of discretization errors, indicating that distinguishing between them might be unnecessary. This message, however, does not hold in the presence of low-dimensional structure, which merits further discussions as follows.
@@
  • Parameterization matters. It is worth noting that how one re-parameterizes the drift term of (5) — namely, how one rearranges the semi-linear drift term therein — can exert significant influences upon the sampling efficiency. This can be elucidated via the inspiring example singled out in Li and Yan, 2024a (, Theorem 2). More concretely, when the target data distribution is isotropic Gaussian on a $k$-dimension linear subspace ($k\ll d$), the iteration complexity of the DDPM sampler is proportional to $O(\mathrm{poly}(k))$ (Li and Yan, 2024a, ), but most of other DDPM-type update rules (i.e., the ones with different coefficients than the original DDPM sampler) can take $\Omega(\mathrm{poly}(d))$ steps to yield the target sampling quality. Informally speaking, parameterizing the semi-linear drift term differently might result in a solution within $O(d/\mathrm{poly}(N))$ KL divergence from the DDPM solution [^3] Roughly speaking, in view of the Taylor expansion, we know that the discrepancy between different solutions is of order $O(1/\mathrm{poly}(N))$; but since $d$-dimensional Gaussian noise is added in each step of the sampling procedure, this discrepancy can result in $O(d/\mathrm{poly}(N))$ difference in KL divergence per step.; this discrepancy is negligible in the full-dimensional setting, but could blow up when $k$ is exceedingly small compared to $d$ and when $N$ is proportional to $k$. All this underscores the critical importance of identifying a parameterization that adapts naturally to the intrinsic data geometry.
@@ 4.1.1 Low dimensionality of the target distribution
    and $\mathsf{dist}(x,\mathcal{M})$ denotes the Euclidean distance of $x$ to $\mathcal{M}$. Also, $\mathsf{vol}(\mathcal{M})$ is the volume, or equivalently, the $k$-dimensional Hausdorff measure, of the manifold $\mathcal{M}$.
    Roughly speaking, the reach of a manifold measures the scale of which the manifold can be regarded as a linear subspace. The larger the reach $\tau_{\mathcal{M}}$ is, the more regular the manifold appears. The volume of the manifold $\mathcal{M}$ is a measure of the size of the manifold, which generalizes the concept of volume in 3-dimensional Euclidean space. Instead of requiring data to be strictly supported on the manifold, Assumption 1 also allows the data to lie in the $\varepsilon_{0}$-enlargement of the manifold $\mathcal{M}$ defined as $\mathcal{M}^{\varepsilon_{0}}:=\{x\in\mathbb{R}^{d}\mid\mathsf{dist}(x,\mathcal{M})\leq\varepsilon_{0}\}$. This is stated formally in the following lemma, with the proof deferred to Appendix C.1.
@@
    If the support of the target data distribution $p_{\mathsf{data}}$ is contained within the manifold $\mathcal{M}^{\varepsilon_{0}}$, i.e., $\mathcal{X}_{\mathsf{data}}\subseteq\mathcal{M}^{\varepsilon_{0}}$,and if $\tau_{\mathcal{M}}=\Omega(k^{-C_{1}})$ and $\mathsf{vol}(\mathcal{M})=O(k^{C_{2}})$ for some universal constants $C_{1},C_{2}>0$, then Assumption 1 holds.
@@
  • Union of the above sets. Our framework can also cover a union of the structures mentioned above, that is, Assumption 1 automatically holds when the data are supported on the union of $k^{O(k)}$ subsets, each of which satisfies Assumption 1.
@@ 4.1.2 Discretization time points
Recall that the discretization time points are $0=t_{0}<t_{1}<\cdots<t_{N}<t_{N+1}=T$, with $T$ the time horizon of the forward process. In addition, we also define throughout the early stopping time
$\delta\coloneqq T-t_{N}<1,$ (14)
the meaning of which shall be discussed momentarily when presenting our convergence theory.
@@
$t_{n+1}-t_{n}\leq\kappa\min\{1,T-t_{n}\}\qquad\text{for all }0\leq n\leq N.$ (15)
@@ 4.1.3 Score estimation errors
$\sum_{n=0}^{N-1}(t_{n+1}-t_{n})\,\mathbb{E}\left[\big{\|}s_{T-t_{n}}(X_{T-t_{n}})-\widehat{s}_{T-t_{n}}(X_{T-t_{n}})\big{\|}_{2}^{2}\right]\leq\varepsilon_{\mathsf{score}}^{2}<\infty,$
@@
Treating the score learning phase as a black box, Assumption 3 singles out a single quantity, $\varepsilon_{\mathsf{score}}\geq 0$, that captures the accuracy of the pre-trained score functions, which is commonly used in recent development of diffusion model theory (Chen et al., 2023a, ; Li et al.,, 2023; Li et al., 2024c, ; Benton et al.,, 2024; Dou et al.,, 2024; Li and Yan, 2024b, ; Li et al., 2024a, ; Li and Yan, 2024a, ; Potaptchik et al.,, 2024; Li and Jiao,, 2024; Azangulov et al.,, 2024; Liang et al.,, 2024). Notably, assuming a weighted average of the mean squared score errors is clearly less stringent than requiring the mean squared score estimation errors to hold uniformly over all time discretization points (Chen et al., 2022b, ; Lee et al.,, 2022; Chen et al., 2023c, ; Wu et al.,, 2024).